PHOTOELECTRON STUDIES OF METHYLNITRENE, CH_{3}N

\underline{BARNEY ELLISON}, Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309-0215.

Complex organic radicals are important intermediates in many combustion and environmental processes. In combustion processes, the organic radical CH is known to cleave N_{2} into HCN and N atoms. Most N atoms are further oxidized to NO which exhausts into the atmosphere. Thus the reaction of CH with N_{2} to produce N atoms connects internal combustion engines with NO production. Some fraction of these N atoms, however, reacts with CH_{3} in the flame to produce CH_{3}N. Methylnitrene is very unstable and decomposes to generate CH_{2}NH or H_{2} + HCN. We will discuss the use of negative ion photoelectron spectroscopy to characterize methylnitrene.

CH_{3}N^{−} + h̄ω_{488 nm} → CH_{3}N + e^{−}

We find EA(CH_{3}N) = 0.022 ± 0.009 eV. In addition to detaching the methylnitrene anion to the ground state of CH_{3}N (X̃ ^{3}A_{2}), we also detect the first electronically excited state of methylnitrene, ã ^{1}E. We measure ΔE(ã ^{1}E−X̃^{3}A_{2}) = 1.352 ± 0.011 eV. The reaction dynamics of nitrene rearrangements will be discussed:

CH_{3}N(X̃ ^{3}A_{2}) → no reaction
CH_{3}N(a^{1}E) → CH_{2}NH